# EF-Calib: Spatiotemporal Calibration of Event- and Frame-Based Cameras Using Continuous-Time Trajectories

### v1

@@ Abstract
Event camera, a bio-inspired asynchronous triggered camera, offers promising prospects for fusion with frame-based cameras owing to its low latency and high dynamic range. However, calibrating stereo vision systems that incorporate both event and frame-based cameras remains a significant challenge. In this letter, we present EF-Calib, a spatiotemporal calibration framework for event- and frame-based cameras using continuous-time trajectories. A novel calibration pattern applicable to both camera types and the corresponding event recognition algorithm is proposed. Leveraging the asynchronous nature of events, a derivable piece-wise B-spline to represent camera pose continuously is introduced, enabling calibration for intrinsic parameters, extrinsic parameters, and time offset, with analytical Jacobians provided. Various experiments are carried out to evaluate the calibration performance of EF-Calib, including calibration experiments for intrinsic parameters, extrinsic parameters, and time offset. Experimental results show that EF-Calib achieves the most accurate intrinsic parameters compared to current SOTA, the close accuracy of the extrinsic parameters compared to the frame-based results, and accurate time offset estimation. EF-Calib provides a convenient and accurate toolbox for calibrating the system that fuses events and frames. The code of this paper will also be open-sourced at: https://github.com/wsakobe/EF-Calib.
@@ I Introduction
Recent years, there has been a growing interest among researchers in a novel bio-inspired camera called the event camera [1]. Abandoning the frame-triggered concept of conventional cameras, each pixel of the event camera can be considered as responding independently and asynchronously to changes in illumination, resulting in an extremely low-latency and high-dynamic-range response pattern. These advantages offer competitive prospects for event cameras in areas such as robotics [2], autonomous driving [3], VR/AR [4], and camera imaging [5].
@@
However, due to the imaging principle of event cameras, they can only react to changes in illumination, making it difficult to capture absolute amounts of illumination and RGB values as frame-based cameras do. This limitation weakens the ability of event cameras to perceive and understand the environment. Therefore, many recent studies have attempted to fuse events with images in order to fully utilize the unique advantages of both modalities, as illustrated in Fig. 1. Some novel SLAM systems achieve more robust localization under fast motion by fusing events and frames [2, 8, 7, 6]. In addition, several studies are exploring how to fuse events and frames for object detection in challenging environments [3, 9]. In recent years, some event-centric datasets with multiple sensors, including frame-based cameras, have also been widely proposed [12, 11, 10].
It is important to note that calibrating the intrinsic and extrinsic parameters of each camera is an indispensable step in the context of multi-camera fusion. Classical camera-to-camera calibration schemes typically require time synchronization, followed by the acquisition of each camera’s parameters through the synchronous acquisition of images of the calibration board with different poses [13]. However, due to the asynchronous nature of the event camera, it is difficult to combine multiple events into single “frames” and time-synchronize them with an image. In addition, events are only generated if there is relative motion, so the event camera cannot capture a stationary calibration board. In summary, a new calibration framework must be designed for the system including event- and frame-based cameras.
To address the aforementioned issues, this letter proposes a novel spatiotemporal calibration framework for event- and frame-based cameras. To the best of our knowledge, EF-Calib is the first calibration framework to achieve joint calibration of event- and frame-based cameras without requiring any time synchronization. The main contributions of this paper are as follows:
@@
    A novel spatiotemporal calibration framework for event- and frame-based cameras is proposed. This framework can jointly obtain the intrinsic and extrinsic parameters, as well as the time offset without requiring any hardware synchronization.
@@ II Related Works
Nevertheless, applying this static and discrete calibration method to event cameras, which are triggered by changes in illumination or relative motion, presents challenges. Initially, many open-source event camera calibration toolkits utilized a synchronized blinking LED calibration board with a known size or a blinking checkerboard pattern generated by an LED screen [14, 15, 16]. This allowed the event camera to identify features similar to a conventional camera and utilize traditional calibration methods. However, these toolkits require complex device preparation and are unsuitable for calibrating extrinsic parameters between event- and frame-based cameras.
In recent years, there has been increased focus on designing new calibration frameworks to facilitate event camera calibration using existing calibration boards. Muglikar et al. [17] utilize deep learning-based image reconstruction networks, such as E2VID [18], to record events generated by moving the calibration board and then apply the reconstructed images to classical calibration methods. However, these methods heavily rely on the quality of image reconstruction and face challenges in achieving time synchronization with conventional cameras. Another approach involves directly utilizing events generated during camera motion for camera calibration. Huang et al. [19] proposed a calibration framework based on a circular calibration board and employed B-splines to optimize the movement trajectory, which is the most similar method to the one proposed in this letter. However, they directly use clustered asynchronous events as features for optimization, compromising sub-pixel accuracy and being highly sensitive to noise. Additionally, their focus is solely on the intrinsic calibration of event cameras, without addressing extrinsic calibration between event cameras and frame-based cameras. Salah et al. [20] also utilize circular calibration boards and introduce eRWLS to fit circular features with sub-pixel accuracy. However, they compress events over a period into a fixed timestamp to obtain a reference “frame” making this method challenging to synchronize with a conventional camera. Furthermore, it does not account for the deformation of circular features at different viewing angles, leading to reduced sub-pixel localization accuracy. The calibration framework proposed in this letter continues this concept and provides an improvement to address the problems of these methods.
@@ III-A Event-Based Vision
Unlike conventional cameras, each pixel of the event camera is independently triggered and responds to changes in the logarithmic illumination signal $L(\mathbf{u}_{k},t_{k})$. An event $(\mathbf{u}_{k},t_{k},p_{k})$ is triggered when the change in logarithmic illumination received by a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$ exceeds a threshold value $C$, i.e.
@@ III-B Continuous-Time Trajectory Representation
Continuous-time trajectories are often represented utilizing a weighted combination of the temporal basis functions [21], such as polynomial functions, FFTs, and Bézier curves. In this letter, the uniform B-spline is introduced as a representation of the continuous-time trajectory. B-splines have the advantages of smoothness, local support, and analytic derivatives, which are well-suited for representing the 6-DoF pose of the event camera [23, 22, 24]. Following the formulation of cumulative $k$th degree B-spline $\mathcal{L}$, the event camera pose $\mathbf{T}^{w}_{e}(\tau)\in\mathbb{SE}\text{3}$ at any time $\tau\in[t_{i},t_{i+1})$, it can be represented by $N$ control points $\mathbf{T}_{i}\in\mathbb{SE}\text{3},i\in[\text{0},\text{1},\ldots,N-\text{1}]$:
$\mathcal{L}:\mathbf{T}_{e}^{w}(\tau)=\mathbf{T}_{i}\cdot\prod_{j=\text{1}}^{k}\mathrm{Exp}\left(\tilde{\mathbf{B}}_{j}(\tau)\cdot\mathrm{Log}\left(\mathbf{T}_{i+j-1}^{-1}\mathbf{T}_{i+j}\right)\right)$ (2)
@@
where $\tilde{\mathbf{M}}^{(k)}\in\mathbb{R}^{(k+\text{1})\times(k+\text{1})}$ is the cumulative blending matrix of B-splines. Since the control points of the B-splines are uniformly distributed on the time scale, the cumulative blending matrix $\mathbf{M}^{(k)}$ is constant. In this letter, considering the continuity and complexity, we use cubic B-splines to represent the camera pose, i.e., $k=\text{3}$. The corresponding cumulative mixing matrix $\tilde{\mathbf{M}}^{(3)}$ [25] is
$\tilde{\mathbf{M}}^{(\text{3})}=\dfrac{\text{1}}{\text{6}}\begin{bmatrix}\text{6}&\text{0}&\text{0}&\text{0}\\ \text{5}&\text{3}&\text{-3}&\text{1}\\ \text{1}&\text{3}&\text{3}&\text{-2}\\ \text{0}&\text{0}&\text{0}&\text{1}\end{bmatrix}$ (5)
Commonly, to simplify the computation, some work decouples the rotation $\mathbf{R}_{e}^{w}(\tau)\in\mathbb{SO}$3 and translation $\mathbf{p}_{e}^{w}(\tau)\in\mathbb{R}^{\text{3}}$ into two independent cubic B-splines, and the same process is carried out in this paper as well. Hence, the continuous-time trajectory of the camera pose can be finally formulated as
$\mathbf{R}_{e}^{w}(\tau)=\mathbf{R}_{i}\cdot\prod_{j=\text{1}}^{3}\mathrm{Exp}\left(\tilde{\mathbf{B}}_{j}(\tau)\cdot\mathrm{Log}\left(\mathbf{R}_{i+j-1}^{-1}\mathbf{R}_{i+j}\right)\right)$ (6)
$\mathbf{p}_{e}^{w}(\tau)=\mathbf{p}_{i}+\sum_{j=\text{1}}^{3}\tilde{\mathbf{B}}_{j}(\tau)\cdot(\mathbf{p}_{i+j}-\mathbf{p}_{i+j-1})$ (7)
After decoupling the pose into two cubic B-splines, the corresponding analytic derivatives [26] can also be derived
$\dot{\mathbf{R}_{e}^{w}}(\tau)$ $=\mathbf{R}_{e}^{w}(\tau)\cdot\left(\boldsymbol{\omega}^{(\text{3})}(\tau)\right)_{\wedge}$ (8)
@@
$\mathbf{v}_{e}^{w}(\tau)=\dot{\mathbf{p}}_{e}^{w}(\tau)=\mathbf{p}_{i}\cdot\sum_{j=\text{1}}^{\text{3}}\dot{\tilde{\mathbf{B}_{j}}}(\tau)\cdot(\mathbf{p}_{i+j}-\mathbf{p}_{i+j-1})$ (9)
where
$\mathbf{A}_{j}=\mathrm{Exp}\left(\tilde{\mathbf{B}}_{j}(\tau)\cdot\mathrm{Log}\left(\mathbf{R}_{i+j-1}^{-1}\mathbf{R}_{i+j}\right)\right)$ (10)
$\dot{\mathbf{A}}_{j}=\mathbf{A}_{j}\dot{\tilde{\mathbf{B}}}(\tau)_{j}\mathrm{Log}\left(\mathbf{R}_{i+j-1}^{-1}\mathbf{R}_{i+j}\right)$ (11)
$\dot{\tilde{\mathbf{B}_{j}}}(\tau)=\dfrac{1}{\Delta t}\tilde{\mathbf{M}}^{(\text{3})}\begin{bmatrix}\text{0}\\[3.00003pt] \text{1}\\[3.00003pt] \text{2}u\\[3.00003pt] \text{3}u^{\text{2}}\end{bmatrix}$ (12)
Since this letter utilizes the uniform B-spline, $\Delta t$ is equal to the time interval between any two consecutive knots, i.e. $\Delta t=t_{i+1}-t_{i}$.
@@ IV-A Calibration Framework
For camera calibration, it is of utmost importance to accurately and robustly identify features on the calibration pattern. However, the checkerboard pattern [27], which is widely used, is difficult to apply to event cameras because events disappear during parallel edge motion. Consequently, calibration patterns for event cameras often employ circular features, but these can produce blur in moving frames. To balance the characteristics of event cameras and frame-based cameras, we designed a new calibration pattern that combines isotropic circles and checkerboard crosspoints, as shown in Fig. 1(a). The center of each circle in this pattern coincides with the center of the inner crosspoint. This hybrid pattern significantly enhances the recognition efficiency and accuracy of the event camera while maintaining compatibility with frame-based cameras.
@@ IV-B Event-Based Feature Recognizer
Unlike frame-based camera, event camera only output asynchronous events during relative motion, posing a challenge for robust calibration pattern recognition. To address this, we propose an event-based calibration pattern feature recognizer, illustrated in Fig. 3. First, we accumulate events over a short period of time $\Delta t$ according to their polarity to obtain “accumulation frames” that resemble traditional images. The introduction of this “accumulation frames” can help us to recognize the feature plate using some classical image processing algorithms. The following subsections describe the recognition algorithm based on “accumulation frames” in detail.
@@ IV-B1 Noise Suppression
Event cameras often generate a large number of events from the static background when they are in motion. These noisy events can adversely affect the recognition of calibration boards, significantly reducing the operation speed and accuracy of the recognizer. Therefore, after obtaining the “accumulation frames”, a noise suppression module is designed to filter out most of the events that are not related to the calibration board.
For circular features, the triggered events typically consist of two semicircular arcs connecting regions of opposite polarity. However, many structures in the background have straight edges, making them more likely to have connected regions that resemble straight lines. To leverage this property, we introduce a fast and accurate connected component labeling (CCL) algorithm called BBDT, proposed by Grana et al. [28]. This algorithm merges neighboring events with the same polarity to obtain all the connectivity regions. Then, the magnitudes of the two principal components of each connected region are calculated using PCA. For background-triggered connected regions, the magnitude of the second principal component $\|\textbf{PC}_{\text{2}}\|$ should be much smaller than the magnitude of the first principal component $\|\textbf{PC}_{\text{1}}\|$, so that a large number of noisy regions can be suppressed by the principal component magnitude ratio $\beta_{PC}$, as given that $\beta_{PC}=\|\textbf{PC}_{\text{1}}\|/\|\textbf{PC}_{\text{2}}\|<T_{PC}$, where $T_{PC}$ is a threshold for the $\beta_{PC}$, and any region with $\beta_{PC}$ higher than $T_{PC}$ is suppressed and not involved in subsequent operations.

### v2

@@ Abstract
The event camera, a bio-inspired asynchronous triggered camera, offers promising prospects for fusion with frame-based cameras owing to its low latency and high dynamic range. However, calibrating stereo vision systems that incorporate both event- and frame-based cameras remains a significant challenge. In this letter, we present EF-Calib, a spatiotemporal calibration framework for event- and frame-based cameras using continuous-time trajectories. A novel calibration pattern applicable to both camera types and the corresponding event recognition algorithm are proposed. Leveraging the asynchronous nature of events, a derivable piece-wise B-spline to represent camera pose continuously is introduced, enabling calibration for intrinsic parameters, extrinsic parameters, and time offset, with analytical Jacobians provided. Various experiments are carried out to evaluate the calibration performance of EF-Calib, including calibration experiments for intrinsic parameters, extrinsic parameters, and time offset. Experimental results demonstrate that EF-Calib outperforms current methods by achieving the most accurate intrinsic parameters, comparable accuracy in extrinsic parameters to frame-based method, and precise time offset estimation. EF-Calib provides a convenient and accurate toolbox for calibrating the system that fuses events and frames. The code of this paper is open-sourced at: https://github.com/wsakobe/EF-Calib.
@@ I Introduction
In recent years, there has been a growing interest among researchers in a novel bio-inspired camera called the event camera [1]. Abandoning the frame-triggered concept of conventional cameras, each pixel of the event camera can be considered as responding independently and asynchronously to changes in illumination, resulting in an extremely low-latency and high-dynamic-range response pattern. These advantages offer competitive prospects for event cameras in areas such as robotics [2], autonomous driving [3, 4], VR/AR [5], and camera imaging [6].
@@
However, due to the imaging principle of event cameras, they only respond to changes in illumination, making it challenging to capture absolute grayscale or RGB values like frame-based cameras. This limitation weakens the ability of event cameras to perceive and understand the environment. Therefore, many recent studies have attempted to fuse events with images in order to fully utilize the unique advantages of both modalities, as illustrated in Fig. 1. Some novel SLAM systems achieve more robust localization under fast motion by fusing events and frames [2, 9, 8, 7]. In addition, several studies are exploring how to fuse events and frames for object detection in challenging environments [3, 10]. In recent years, some event-centric datasets with multiple sensors, including frame-based cameras, have also been widely proposed [13, 12, 11].
It is important to note that calibrating the intrinsic and extrinsic parameters of each camera is an indispensable step in the context of multi-camera fusion. Classical multi-camera calibration frameworks typically require time synchronization among all cameras, followed by the acquisition of each camera’s parameters through the synchronous capture of images of a calibration board in various poses [14]. However, due to the asynchronous nature of event cameras, it is challenging to aggregate multiple events into single “frames” and time-synchronize them with traditional images. Additionally, events are generated only when the illumination at each pixel changes. As a result, a static event camera cannot capture static objects without any illumination change. In conclusion, a user-friendly, and cost-effective calibration framework is needed for systems that include both event- and frame-based cameras.
To address the aforementioned issues, this letter proposes a novel spatiotemporal calibration framework for event- and frame-based cameras called EF-Calib. To the best of our knowledge, EF-Calib is the first calibration framework to achieve joint calibration of event- and frame-based cameras without requiring any time synchronization. The main contributions of this paper are as follows:
@@
    A novel spatiotemporal calibration framework for event- and frame-based cameras is proposed. This framework can obtain the intrinsic and extrinsic parameters, as well as the time offset without requiring any hardware synchronization.
@@ II Related Works
Nevertheless, applying this static and discrete calibration method to event cameras, which are triggered by changes in illumination or relative motion, presents challenges. Initially, many open-source event camera calibration kits utilized a synchronized blinking LED calibration board [15, 16, 17], enabling alignment between cumulative event camera frames and traditional images. However, this approach requires complex hardware, leading to higher costs and reduced calibration accuracy. Alternatively, static calibration can be achieved by adjusting external light intensity, but this introduces significant noise and fails to account for time offset between cameras. As a result, these toolkits are not the most efficient solution for calibrating event-based and frame-based cameras.
In recent years, there has been increased focus on designing new calibration frameworks to facilitate event camera calibration using existing calibration boards. Muglikar et al. [18] utilize deep learning-based image reconstruction networks, such as E2VID [19], to record events generated by moving the calibration board and then apply the reconstructed images to classical calibration methods. However, these methods heavily rely on the quality of image reconstruction and face challenges in achieving time synchronization with conventional cameras. Another approach involves directly utilizing events generated during camera motion for camera calibration. Huang et al. [20] proposed a calibration framework based on a circular calibration board and employed B-splines to optimize the movement trajectory, which is the most similar method to the one proposed in this letter. However, they directly use clustered asynchronous events as features for optimization, compromising sub-pixel accuracy and being highly sensitive to noise. Salah et al. [21] also utilize circular calibration boards and introduce eRWLS to fit circular features with sub-pixel accuracy. However, they compress events over a period into a fixed timestamp to obtain a reference “frame” making this method challenging to synchronize with a frame-based camera. Furthermore, it does not account for the deformation of circular features at different viewing angles, leading to reduced sub-pixel localization accuracy. The calibration framework proposed in this letter continues this concept and provides an improvement to address the problems of these methods.
@@ III-A Event-Based Vision
Unlike conventional cameras, each pixel of the event camera is independently triggered and responds to changes in the logarithmic illumination signal $L(\mathbf{u}_{k},t_{k})$. An event $(\mathbf{u}_{k},t_{k},p_{k})$ is triggered when the change in logarithmic illumination received by a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$ exceeds a threshold value $C$, i.e.,
@@ III-B Continuous-Time Trajectory Representation
Continuous-time trajectories are often represented utilizing a weighted combination of the temporal basis functions [22], such as polynomial functions, FFTs, and Bézier curves. In this letter, the uniform B-spline is introduced as a representation of the continuous-time trajectory. B-splines have the advantages of smoothness, local support, and analytic derivatives, which are well-suited for representing the 6-DoF pose of the event camera [24, 23, 25]. Following the formulation of cumulative $k$th degree B-spline $\mathcal{L}$, the event camera pose $\mathbf{T}^{e}_{w}(\tau)\in\mathbb{SE}\text{3}$ at any time $\tau\in[t_{i},t_{i+1})$, it can be represented by $N$ control points $\mathbf{T}_{i}\in\mathbb{SE}\text{3},i\in[\text{0},\text{1},\ldots,N-\text{1}]$:
$\mathcal{L}:\mathbf{T}_{w}^{e}(\tau)=\mathbf{T}_{i}\cdot\prod_{j=\text{1}}^{k}\mathrm{Exp}\left(\tilde{\mathbf{B}}_{j}(\tau)\cdot\mathrm{Log}\left(\mathbf{T}_{i+j-1}^{-1}\mathbf{T}_{i+j}\right)\right)$ (2)
@@
where $\tilde{\mathbf{M}}^{(k)}\in\mathbb{R}^{(k+\text{1})\times(k+\text{1})}$ is the cumulative blending matrix of B-splines. Since the control points of the B-splines are uniformly distributed on the time scale, the cumulative blending matrix $\mathbf{M}^{(k)}$ is constant. In this letter, considering the continuity and complexity, we use cubic B-splines to represent the camera pose, i.e., $k=\text{3}$. The corresponding cumulative mixing matrix $\tilde{\mathbf{M}}^{(3)}$ can be calculated as described in [26].
Representing B-splines using (2) leads to computational inefficiency and coupling issues. According to [27, 28], decoupling $\mathbf{T}_{w}^{e}(\tau)\in\mathbb{SE}$3 into rotations $\mathbf{R}_{w}^{e}(\tau)\in\mathbb{SO}$3 and translations $\mathbf{t}_{w}^{e}(\tau)\in\mathbb{R}^{\text{3}}$ improves both motion trajectory representation and computational time. Hence, the continuous-time trajectory of the camera pose can be finally formulated as
$\mathbf{R}_{w}^{e}(\tau)=\mathbf{R}_{i}\cdot\prod_{j=\text{1}}^{3}\mathrm{Exp}\left(\tilde{\mathbf{B}}_{j}(\tau)\cdot\mathrm{Log}\left(\mathbf{R}_{i+j-1}^{-1}\mathbf{R}_{i+j}\right)\right)$ (5)
$\mathbf{t}_{w}^{e}(\tau)=\mathbf{t}_{i}+\sum_{j=\text{1}}^{3}\tilde{\mathbf{B}}_{j}(\tau)\cdot(\mathbf{t}_{i+j}-\mathbf{t}_{i+j-1})$ (6)
After decoupling the pose trajectory into two cubic B-splines, the corresponding analytic derivatives [27] can also be derived
$\dot{\mathbf{R}_{w}^{e}}(\tau)$ $=\mathbf{R}_{w}^{e}(\tau)\cdot\left(\boldsymbol{\omega}^{(\text{3})}(\tau)\right)_{\wedge}$ (7)
@@
$\mathbf{v}_{e}^{w}(\tau)=\dot{\mathbf{t}}_{e}^{w}(\tau)=\mathbf{t}_{i}\cdot\sum_{j=\text{1}}^{\text{3}}\dot{\tilde{\mathbf{B}_{j}}}(\tau)\cdot(\mathbf{t}_{i+j}-\mathbf{t}_{i+j-1})$ (8)
where $\mathbf{A}$, $\dot{\mathbf{A}}$, and $\dot{\tilde{\mathbf{B}}}$ can also be found in [27]. Since this letter utilizes the uniform B-spline, $\Delta t$ is equal to the time interval between any two consecutive knots, i.e., $\Delta t=t_{i+1}-t_{i}$.
@@ IV-A Calibration Framework
For camera calibration, it is of utmost importance to accurately and robustly identify features on the calibration pattern. However, the checkerboard pattern [29], which is widely used, is difficult to apply to event cameras because events disappear during parallel edge motion [1]. Consequently, calibration of event cameras typically uses circular features. However, for moving frame-based cameras, circular features are more severely affected by motion blur than checkerboard crosspoints. To balance the needs of both event cameras and frame-based cameras, we designed a new calibration pattern that combines isotropic circles with checkerboard crosspoints, as illustrated in Fig. 1(a). The center of each circle in this pattern coincides with the center of the inner crosspoint. This hybrid pattern significantly enhances the recognition efficiency and accuracy of the event camera while maintaining compatibility with frame-based cameras.
@@ IV-B Event-Based Feature Recognizer
Unlike frame-based cameras, event cameras only output asynchronous events when the illumination of each pixel changes over a threshold, posing a challenge for robust calibration pattern recognition. To address this, we propose an event-based calibration pattern feature recognizer, illustrated in Fig. 3. First, we accumulate events over a short period of time $\Delta t$ according to their polarity to obtain “accumulation frames” that resemble traditional images. The introduction of this “accumulation frames” can help us to recognize the feature plate using some classical image processing algorithms. The following subsections describe the recognition algorithm based on “accumulation frames” in detail.
@@ IV-B1 Noise Suppression
Event cameras typically generate a large number of events from a static background when in motion. These noisy events do not contribute to the recognition of the calibration board and instead significantly increase the computational load on the feature extraction module, thereby reducing the efficiency of the calibration process. To address this, after obtaining the “accumulation frames”, a noise suppression module should be implemented to filter out most of the events that are unrelated to the calibration board.
For circular features, the triggered events typically consist of two semicircular arcs connecting regions of opposite polarity. However, many structures in the background have straight edges, making them more likely to have connected regions that resemble straight lines. To leverage this property, we introduce a fast and accurate connected component labeling (CCL) algorithm called BBDT [30]. This algorithm merges neighboring events with the same polarity to obtain all the connectivity regions. Then, the magnitudes of the two principal components of each connected region are calculated using PCA. For background-triggered connected regions, the magnitude of the second principal component $\|\textbf{PC}_{\text{2}}\|$ should be much smaller than the magnitude of the first principal component $\|\textbf{PC}_{\text{1}}\|$, so that a large number of noisy regions can be suppressed by the principal component magnitude ratio $\beta_{PC}$, as given that $\beta_{PC}=\|\textbf{PC}_{\text{1}}\|/\|\textbf{PC}_{\text{2}}\|<T_{PC}$, where $T_{PC}$ is a threshold for the $\beta_{PC}$, and any region with $\beta_{PC}$ higher than $T_{PC}$ is suppressed and not involved in subsequent operations.